OpenStudy (lgbasallote):

@sami-21 can i use \[\mathcal L \lbrace t^n f(t) \rbrace = (-1)^n \frac{d^n}{ds^n} \left[f(s)\right]\]in \[\mathcal L \lbrace t^{1/2} \rbrace\]

OpenStudy (anonymous):

nooooopes ! it is something different

OpenStudy (lgbasallote):

i know for sure i cant do \[\large \frac{n!}{s^{n+1}}\]

OpenStudy (lgbasallote):

right?

OpenStudy (anonymous):

yes you cannot use that as well !!

OpenStudy (anonymous):

here gamma function values is required

OpenStudy (lgbasallote):

uhh no we didnt have that

OpenStudy (lgbasallote):

but fine...teach it to me

OpenStudy (anonymous):

for such powers of t you need \[\Large L(t^n)=\frac{\Gamma(n+1)}{s^{n+1}}\]

OpenStudy (anonymous):

@mukushla am i right ?

OpenStudy (lgbasallote):

what do these mean o.O

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

ok using n=1/2 we have in the above formula \[\Large L(t^{1/2})=\frac{\Gamma(1/2+1)}{s^{1/2+1}}=\frac{\Gamma(3/2)}{s^{3/2}}\]

OpenStudy (anonymous):

where \[\Large \Gamma(3/2)=\frac{\sqrt{\pi}}{2}\]

OpenStudy (lgbasallote):

wanna explain what that inverted L is and how to do it in latex?

OpenStudy (lgbasallote):

WOAHHHH`

OpenStudy (anonymous):

it is not possible for me to explain here in detail gamma function .but i can tell you that much so that you will easily be able to complete your Laplace course. for this you have to remember the following \[\Large \Gamma(1/2)=\sqrt{\pi}\] just remember the following formula \[\Large \Gamma(n+1)=n \Gamma(n)\] let me explain below.

OpenStudy (lgbasallote):

how to do that in latex? \(\gamma\)?

OpenStudy (lgbasallote):

\[\Gamma\]

OpenStudy (lgbasallote):

ahh found it

OpenStudy (anonymous):

so if you need gamma(1/2+1) here n=1/2 using the above formula nd known value of gamma(1/2)=sqrt(pi} \[\Large \Gamma(3/2)=\Gamma(1/2+1)=1/2\Gamma(1/2)=1/2\sqrt{\pi}\] similarly for Gamma(3/2+1) here n=3/2 using above value of gamma(3/2 ) we have \[\Large \Gamma(5/2)=\frac{3}{2}*\frac{\sqrt{\pi}}{2}\] can you now tell me \[\Large \Gamma(\frac{7}{2})\] ??

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

this gamma thing...when is it used again??

OpenStudy (anonymous):

for example \[\Huge L(t^{-1/2})\] you will need this gamma thing

OpenStudy (unklerhaukus):

\[\Gamma (n)=(n-1)!\]

OpenStudy (lgbasallote):

so basially when the given is \[\large \mathcal L \lbrace t^n \rbrace\] yes?

OpenStudy (anonymous):

@UnkleRhaukus yes you are right , but it will not be helpful if you have n=1/2 ,3/2,5/2...

OpenStudy (anonymous):

you could use: \[L(t^{n-1/2})=((1*3*5*...*(2n-1))\sqrt \pi)/(2^ns^{n+1/2})\]

OpenStudy (anonymous):

where n is natural number

OpenStudy (lgbasallote):

let's stick to one methid to avoid confusing teh asker shall we.... :(

OpenStudy (anonymous):

so just set n=1, to get laplace transform of t^1/2

OpenStudy (unklerhaukus):

then you'd have to use \[\Gamma(z+1)=z\Gamma(z)\] & \[\Gamma \left(\frac12 \right)=\sqrt{\pi}\]

OpenStudy (lgbasallote):

so this gamma thing is used whenever the given is L{t^n}???

OpenStudy (anonymous):

i do not know my screen automatically becomes blank :(

OpenStudy (lgbasallote):

refresh?

OpenStudy (anonymous):

ok just to write that all again was vanished by blanking of screen:( here it is when n=0,1,2,3,4,5... then \[\Large \Gamma(n+1)=n!\] and you r known formula for n=0,1,2,3.. is used \[\Large L(t^n)=\frac{n!}{s^{n+1}}\]

OpenStudy (unklerhaukus):

This is a standard result in any laplace transform table \[ \mathcal L\{t^r\}=\frac{\Gamma(r+1)}{s^{r+1}}\qquad; ~r>-1\]

OpenStudy (anonymous):

but if n is n0t 0,1,2,3,4.......then you will have to use \[\Large f(t^n)=\frac{\Gamma(n+1)}{s^{n+1}}\]

OpenStudy (unklerhaukus):

when r is a natural number n; \(\Gamma(n+1)=n!\)

OpenStudy (lgbasallote):

WOAH...what's happening here??!! o.O

OpenStudy (lgbasallote):

i was still asking a question...

OpenStudy (lgbasallote):

is this gamma thingy done when the given is L{t^n}??

OpenStudy (unklerhaukus):

if n is a natural number you might as well go straight to n factorial but if is is half integer use the gamma function and \(\Gamma(1/2)=\sqrt\pi\)

OpenStudy (lgbasallote):

okay....in sami's earlier solution i was able to understand as far as \[\frac 12 \sqrt \pi\] after that im lost

OpenStudy (unklerhaukus):

what bit dont you get? @lgbasallote ?

OpenStudy (lgbasallote):

similarly for Gamma(3/2+1) here n=3/2 using above value of gamma(3/2 ) we have \[\Gamma (\frac 52) = \frac 32 \times \frac{\sqrt{\pi}}{2}\] and so on...

OpenStudy (lgbasallote):

look at sami's earlier post...

OpenStudy (lgbasallote):

ing to that one

OpenStudy (lgbasallote):

i dont get what he did after \[\frac 12 \Gamma \frac 12 \implies \frac{\sqrt{\pi}}{2}\]

OpenStudy (lgbasallote):

oh wait....he was just giving examples....stupid me...

OpenStudy (lgbasallote):

i thought it was related to the problem

OpenStudy (lgbasallote):

so how do i use the fact that \[\Gamma (1/2) = \sqrt \pi\] and \[\Gamma (n+1) = n\Gamma (n)\] to solve \[\mathcal L \lbrace t^{1/2} \rbrace\]

OpenStudy (anonymous):

Read this http://en.wikipedia.org/wiki/Gamma_function

OpenStudy (unklerhaukus):

\[\Gamma\left(\frac{7}{2}\right)=\frac{5}{2}\Gamma\left(\frac{5}{2}\right)=\frac{5}{2}\cdot\frac{3}{2}\Gamma\left(\frac{3}{2}\right)=\frac{5}{2}\cdot\frac{3}{2}\cdot\frac12\Gamma\left(\frac{1}{2}\right)=\frac{5}{2}\cdot\frac{3}{2}\cdot\frac12\cdot \sqrt{\pi}\]

OpenStudy (lgbasallote):

wait i think i get it now... sami said \[f(t^n) = \frac{\Gamma(n+1)}{s^{n+1}}\] so \[\mathcal L \lbrace t^{1/2} \rbrace = \frac{\Gamma(\frac 12 + 1)}{s^{1/2 + 1}}\] right?

OpenStudy (lgbasallote):

is that right @sami-21 ? because if yes...then i think im getting it now

OpenStudy (unklerhaukus):

and , simplify ..

OpenStudy (lgbasallote):

nice.. \[\frac{1}{2s} \left(\frac{\pi}{s}\right)^{1/2} \quad s> 0\] just like what the book says

OpenStudy (anonymous):

@lgbasallote it is correct !!

OpenStudy (anonymous):

@lgbasallote yes you are a fast learner :)

OpenStudy (lgbasallote):

lol on contraire...it took me an hour to get what you said =)))

OpenStudy (lgbasallote):

wow..are you a mathematician sir? your solutions are somehow...mathematical...idk hoq to explain it

OpenStudy (anonymous):

That is the theory in general: \[ \Gamma(z)=\int_0^\infty e^{-t} t^{z-1} dt\\ \cal L( t^{n})=\int_0^\infty e^{-s t} t^{n} dt\\ u= s t,\, du = s dt, \, dt =\frac {du}{ s}, \, t =\frac u s\\ \cal L( t^{n})=\int_0^\infty e^{-u} (u/s)^{n}\frac {du}{s}=\\ \frac 1 {s^{3/2}}\int_0^\infty e^{-u} u^{n}du=\\ \frac 1 {s^{n+1}}\int_0^\infty e^{-u} u^{n}du=\frac {\Gamma(n+1)}{s^{n+1}} \]

OpenStudy (lgbasallote):

is this the proof?

OpenStudy (anonymous):

That is the theory in general: \[ \Gamma(z)=\int_0^\infty e^{-t} t^{z-1} dt\\ \cal L( t^{n})=\int_0^\infty e^{-s t} t^{n} dt\\ u= s t,\, du = s dt, \, dt =\frac {du}{ s}, \, t =\frac u s\\ \cal L( t^{n})=\int_0^\infty e^{-u} (u/s)^{n}\frac {du}{s}=\\ \frac 1 {s^{n+1}}\int_0^\infty e^{-u} u^{n}du=\\ \frac 1 {s^{n+1}}\int_0^\infty e^{-u} u^{n}du=\frac {\Gamma(n+1)}{s^{n+1}} \]

OpenStudy (anonymous):

That is the proof in general. In your solution, you can use this fact without proving it.

OpenStudy (lgbasallote):

i just wonder though...how did \[\Gamma (n+1) \] become n! if n is an integer?

OpenStudy (anonymous):

The \(\Gamma\) funtcion is a generalization of the factorial function.

OpenStudy (anonymous):

Look at the definition of the \(\Gamma \) function. (First line in my post)

OpenStudy (lgbasallote):

ohh...nevermind i found out.. like for example n = 3 \[\Gamma (3+1) = 3\Gamma (3) \implies 3 \Gamma (2+1) \implies 3(2) \Gamma (1)\] right?

OpenStudy (anonymous):

Yes. \(\Gamma(1)=1\)

OpenStudy (lgbasallote):

nice. thanks. that makes sense

OpenStudy (anonymous):

In general \( \Gamma(n+1) = n! \) if n is a postive integer.

OpenStudy (anonymous):

Try to prove that for any n integer or not \[ \Gamma(n+1) = n \Gamma(n) \]

OpenStudy (lgbasallote):

n cannot be less than 0 yes?

OpenStudy (anonymous):

Yes.

OpenStudy (lgbasallote):

okay..let me try to prove that...

OpenStudy (anonymous):

Yes if n is a negative integer, but n can be for example -1/3

OpenStudy (lgbasallote):

ohh it can be negative fraction but not negative integer...makes sense

OpenStudy (anonymous):

Since I have to go for my hike \[ \Gamma(z)=\int_0^\infty e^{-t} t^{z-1} dt\\ \Gamma(n+1)=\int_0^\infty e^{-t} t^{n} dt\\ u=t^n,\, dv = e^{-t} dt\\ du= nt^{n-1} dt, \, v=-e^{-t}\\ \Gamma(n+1)=\int_0^\infty e^{-t} t^{n} dt=\\ \left[- t^n e^{-t} \right]_{t=0}^\infty+ n \int_0^\infty e^{-t} t^{n-1} dt=0+ n\Gamma(n) \] I used integration by parts above.

OpenStudy (anonymous):

@eliassaab thanks .you really made it clear to him. @lgbasallote hope you got it now ! :) Best of Luck with Laplace :)

OpenStudy (anonymous):

yw